# Rotatable Antenna-array-enhanced Direction-sensing for Low-altitude Communication Network: Method and Performance

### v1

@@ Abstract
In a practical multi-antenna receiver, each element of the receive antenna array has a directive antenna pattern, which is still not fully explored and investigated in academia and industry until now. When the emitter is deviated greatly from the normal direction of antenna element or is close to the null-point direction, the sensing energy by array will be seriously attenuated such that the direction-sensing performance is degraded significantly. To address such an issue, a rotatable array system is established with the directive antenna pattern of each element taken into account, where each element has the same antenna pattern. Then, the corresponding the Cramer-Rao lower bound (CRLB) is derived. Finally, a recursive rotation Root-MUSIC (RR-Root-MUSIC) direction-sensing method is proposed and its mean-square-error (MSE) performance is evaluated by the derived CRLB. Simulation results show that the proposed rotation method converges rapidly with about ten iterations, and make a significant enhancement on the direction-sensing accuracy in terms of MSE when the target direction departs seriously far away from the normal vector of array. Compared with conventional Root-MUSIC, the sensing performance of the proposed RR-Root-MUSIC method is much closer to the CRLB.
@@ I Introduction
    To overcome the serious receive energy attenuation caused by the antenna element directive pattern, a rotatable planar array system is established to enhance the sensing performance where each element is assumed to have the same antenna pattern. Different from the conventional fixed array system, the proposed system may near the emitter direction by multiple operations of sensing and rotation to improve the receive signal-to-noise ratio (SNR). As SNR grows, the sensing accuracy will be improved correspondingly. The multiple sensing and rotation framework will be very suitable for the future low-altitude wireless networks with rich UAVs.
@@
    Simulation is conducted to assess the error performance of the proposed RR-Root-MUSIC. In accordance with simulation results, the proposed RR-Root-MUSIC performs much better than conventional fixed Root-MUSIC and is closer to the corresponding CRLB. More important, the proposed method has a fast convergent speed of about ten iterations. When the target direction departs far away from the normal vector of array, the performance gain achieved by recursive sensing and rotation process is tremendous. For example, at $\theta\leq 15^{\circ}$, where $\theta$ denotes the direction difference between the target direction and the antenna array plane, the achievable performance gain of the proposed RR-Root-MUSIC over conventional Root-MUSIC will be up to about six-order magnitude or more. In particular, as $\theta$ nears the array plane, i.e. far away from the normal direction of rotation array, the achievable performance gain will grow gradually and become more significant.
@@ II System Model
Let us define that $\mathbf{y}\in\mathbb{C}^{1\times NM}$ and $\mathbf{n}\in\mathbb{C}^{1\times NM}$ represent the received data matrix and the AWGN noise matrix, respectively. Assume the mean value of AWGN noise is zero and the mean-square-error (MSE) is $\sigma^{2}$. Then, $\mathbf{y}$ can be repressed as
@@
$\mathbf{a}_{x}(\theta,\phi)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}x_{0}\sin\theta\cos\phi}\\ \cdots\\ e^{j\frac{2\pi}{\lambda}x_{n}\sin\theta\cos\phi}\\ \cdots\\ e^{j\frac{2\pi}{\lambda}x_{N}\sin\theta\cos\phi}\end{bmatrix}$, (5)
@@
$\mathbf{a}_{z}(\theta)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}z_{0}\cos\theta}\cdots e^{j\frac{2\pi}{\lambda}z_{m}\cos\theta}\cdots e^{j\frac{2\pi}{\lambda}z_{M}\cos\theta}\end{bmatrix}^{H}.$ (6)
@@
$\tilde{\mathbf{a}}_{z}(\theta)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}x_{0}\sin\theta\cos(\phi+\Delta_{\phi})}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}x_{n}\sin\theta\cos(\phi+\Delta_{\phi})}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}x_{N}\sin\theta\cos(\phi+\Delta_{\phi})}\end{bmatrix}$, (16)
@@
$\tilde{\mathbf{a}}_{x}(\theta,\phi)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}z_{0}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}z_{m}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}z_{M}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\end{bmatrix}$. (17)
@@ III Proposed Recursive Rotation Root-MUSIC (RR-Root-MUSIC) Method
Due to directive antenna patterns, the sensing power is significantly attenuated when the emitter direction deviates from the normal vector. Therefore, in this section, a RR-Root-MUSIC method is proposed. As shown in Fig. 3LABEL:sub@a, when the emitter direction is far away from the normal vector of array, Root-MUSIC method is applied at the initial orientation which provides a coarse estimation value at the lower directional gain. According to this estimation angle, the array is rotated into the main-lobe region as Fig. 3LABEL:sub@b, which increases the sensing power and improves the estimation condition. With the updated orientation, a new estimation angle is obtained, and the process repeats. As shown in Fig. 3LABEL:sub@c, the offset angle gradually shrinks over iterations, leading to higher effective SNR and eventual convergence toward the lower MSE.
@@ IV Derived CRLB for Rotatable Array
The CRLB provides a benchmark for the minimum achievable MSE of any unbiased estimator. According to literature [34], [35] and [36], to evaluate the performance of the proposed RR-Root-MUSIC method, the CRLB of the rotation array is derived. The derivation process is as follows.
@@
$\tilde{\mathbf{y}}=\big[\tilde{y}_{0,0},\,\tilde{y}_{0,1},\dots,\tilde{y}_{M-1,N-1}\big]^{T},$
@@
$\boldsymbol{\theta}=[\theta,\phi]^{T}.$
@@
$\mathbf{F}(\theta,\varphi)=\frac{2\,K\,s^{2}}{\sigma^{2}}\begin{bmatrix}P&Q\\ Q&R\end{bmatrix}$. (47)
@@
$CRLB=\mathbf{F}^{-1}(\theta,\varphi)=\frac{\sigma^{2}}{2Ks^{2}}\frac{1}{PR-Q^{2}}\begin{bmatrix}P&-Q\\ -Q&R\end{bmatrix}$. (57)
Therefore, the lower bound of the MSE of the $\hat{\theta}$ and $\hat{\phi}$ can be expressed as
@@ V Simulation Results
In this section, we present simulation results to evaluate the performance of proposed RR-Root-MUSIC method and the corresponding CRLB as a performance benchmark. The parameters of simulation are set as: $N=M=6$, $P_{t}$= 20dBm, $\sigma^{2}$= -100dBm, $\phi=90^{\circ}$, $r$ = 250m, $\lambda$ = 0.125. In our simulation, all results are averaged over 2000 Monte Carlo realizations and MSE is calculated to indicate the performance, which is given by
$\text{MSE}=\frac{1}{L}\sum_{l}^{L}(\theta_{l}-\theta)^{2},$ (66)
@@
In Fig. 4, the MSE convergence curves of the proposed RR-Root-MUSIC method are illustrated for SNR = -10 dB and three different elevation angles: $\theta=75^{\circ}$, $\theta=45^{\circ}$, $\theta=15^{\circ}$, with CRLB as a performance benchmark, where $L=1$ in Fig. 4LABEL:sub@fig:conver-curve-low_L1 and $L=1000$ in in Fig. 4LABEL:sub@fig:conver-curve-low_L1000. Obviously, given a fixed value of SNR, the proposed method asymptotically converges to the corresponding error floor after about 13 or less rotations on average. As the angle of the emitter deviates away from the normal vector of the rotation array, the MSE of the harvested performance gain ranges from one-order magnitude to six-order magnitude as $\theta$ varies from $75^{\circ}$ to $15^{\circ}$. This means that the larger the angle difference between the incident direction of the emitter and the normal direction of array, the larger the achieved performance gain by the proposed rotation operation. When $\theta$ is close to $90^{\circ}$, the proposed method takes only one time to converge to the same error floor. In summary, if the emitter direction is far away from the normal vector, a significant MSE performance gain may be achieved.
@@
Figure 7: MSE versus the elevation angles $\theta$ for proposed RR-Root-MUSIC method with transmit powers $P_{t}\in$ {0 dBm, 20 dBm}.
In Fig. 7, the MSE curves of the proposed RR-Root-MUSIC method versus the elevation angle $\theta$ are shown for different transmitter powers $P_{t}\in$ {0 dBm, 20 dBm}, with CRLB as a performance benchmark. It is obvious that as $\theta\rightarrow 0^{o}$, the performance gain of the proposed RR-Root-MUSIC over conventional fixed Root-MUSIC grows gradually. In extreme situation, i.e., $\theta\leq 5^{o}$, the MSE performance of the proposed method is more eight-order magnitude better than that of the conventional fixed Root-MUSIC. This further confirms that a rotatable array may effectively mitigate the pattern-induced performance loss.
@@
(b) MSE versus the motion path of UAV
Figure 8: The motion path of the UAV around BS, and MSE of proposed RR-Root-MUSIC method versus the motion path of UAV.
To observe the influence of the proposed rotation method in low-latitude UAV flying, Fig. 8LABEL:sub@fig:uav_path shows the motion path of the UAV around BS. Fig. 8LABEL:sub@fig:MSE_uav illustrates the corresponding MSE performance curves along semi-circular trajectory of UAV for the proposed RR-Root-MUSIC and the traditional fixed MUSIC method. From Fig. 8LABEL:sub@fig:uav_path, the UAV flies form (0, -250, 0) and (0, -150, 0) to (0, 250, 0) and (0, 150, 0) with radii $r=250$ and $r=150$, respectively. Compared with the traditional Root-MUSIC method, the proposed RR-Root-MUSIC method makes a dramatic performance gain over the conventional fixed Root-MUSIC when the position of the UAV is on top of BS, which is usually a blind coverage area. Additionally, due to the use of rotatable array, the sensing performance keeps stable during the flight process of UAV. Thus, the proposed rotatable array system and method will offer an excellent new solution to the future low-latitude communication network.
Figure 9: MSE of proposed RR-Root-MUSIC method versus the boresight deflection angle $\varphi$ for the proposed method with the directivity factor $p\in$ {0, 0.5, 1, 2}.
Fig. 9 plots the versus between MSE and the boresight deflection angle $\varphi$ for the proposed RR-Root-MUSIC method under directivity factors $p\in$ {0, 0.5, 1, 2}. When the elevation and azimuth angle are fixed and the boresight deflection angle varies, it is observed that as $\varphi$ increases from 0 to $\frac{\pi}{2}$ or decreases from 0 to $\frac{\pi}{2}$, the MSE decreases significantly. This phenomenon is attributed to the directive antenna pattern. At large boresight deflection angle, the directional gain is severely attenuated, leading to a reduced SNR. The MSE is more sensitive to $\varphi$, yielding a larger gain in the main lobe and greater gain attenuation in directions deviating from the boresight. Compared with an isotropic array, the directive array achieves better performance near the boresight direction.
@@ VI Conclusions
In this paper, to solve the low-altitude coverage problem of MIMO transceiver with directional antenna elements in practical, a rotatable array system was established with each element being the same directional pattern. Here, via a sensing and rotation iterative loop, the RR-Root-MUSIC was proposed to fully harvest the antenna array gain to make a dramatic performance enhancement, especially for the scenario that the emitter direction is far away from the normal vector of array. Compared with classical Root-MUSIC without rotation, the proposed method may achieve up to six-order magnitude improvement or more at the elevation angle $\theta\leq 15^{o}$ in terms of MSE performance. Moreover, the proposed RR-Root-MUSIC converges fast and provides a full coverage on such an area on top of BS, which cannot be covered by the conventional fixed antenna array system like 5G or 4G. Due to the above advantages of rotation-array system, it will become an extremely potential technique for the future low-altitude wireless network to completely break the bottleneck of low-latitude coverage faced by the current mobile communication like 5G.

### v2

@@ Abstract
In a practical multi-antenna receiver, each element of the receive antenna array has a directive antenna pattern, which is still not fully explored and investigated in academia and industry until now. When the emitter is deviated greatly from the normal direction of antenna element or is close to the null-point direction, the sensing energy by array will be seriously attenuated such that the direction-sensing performance is degraded significantly, which will be faced in the low-altitude wireless network with a large amount of unmanned aerial vehicles (UAVs) in the coming future. To address such an issue, a rotatable array system is established with the directive antenna pattern of each element taken into account, where each element has the same antenna pattern. Then, the corresponding the Cramer-Rao lower bound (CRLB) is derived. Finally, a recursive rotation Root-MUSIC (RR-Root-MUSIC) direction-sensing method is proposed and its root-mean-squared error (RMSE) performance is evaluated by the derived CRLB. Simulation results show that the proposed rotation method converges rapidly with about ten iterations, and make a significant enhancement on the direction-sensing accuracy in terms of RMSE when the target direction departs seriously far away from the normal vector of array. Compared with conventional Root-MUSIC, the sensing performance of the proposed RR-Root-MUSIC method is much closer to the CRLB.
@@ I Introduction
    To overcome the serious receive energy attenuation caused by the antenna element directive pattern, a rotatable planar array system is established to enhance the sensing performance where each element is assumed to have the same antenna pattern. Different from the conventional fixed array system, the proposed system may near the emitter direction by multiple operations of sensing and rotation to improve the receive SNR. As SNR grows, the sensing accuracy will be improved correspondingly. The multiple sensing and rotation framework will be very suitable for the future low-altitude wireless networks with rich UAVs.
@@
    Simulation is conducted to assess the error performance of the proposed RR-Root-MUSIC. In accordance with simulation results, the proposed RR-Root-MUSIC performs much better than conventional fixed Root-MUSIC and is closer to the corresponding CRLB. More important, the proposed method has a fast convergent speed of about ten iterations. When the target direction departs far away from the normal vector of array, the performance gain achieved by recursive sensing and rotation process is tremendous. For example, at $\theta\leq 15^{\circ}$, where $\theta$ denotes the direction difference between the target direction and the antenna array plane, the achievable performance gain of the proposed RR-Root-MUSIC over conventional Root-MUSIC will be up to about three-order magnitude or more. In particular, as $\theta$ nears the array plane, i.e. far away from the normal direction of rotation array, the achievable performance gain will grow gradually and become more significant.
The remainder of this research is organized as follows. The system model of rotatable array is described in Section II. In Section III, the RR-Root-MUSIC is proposed. In Section IV, corresponding CRLB is also derived. Moreover, section V presents the experimental results, with conclusions provided in Section VII.
@@ II System Model
Let us define that $\mathbf{y}\in\mathbb{C}^{1\times NM}$ and $\mathbf{n}\in\mathbb{C}^{1\times NM}$ represent the received data matrix and the AWGN noise matrix, respectively. Assume the mean value of AWGN noise is zero and the root-mean-squared error (RMSE) is $\sigma^{2}$. Then, $\mathbf{y}$ can be repressed as
@@
$\mathbf{a}_{x}(\theta,\phi)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}x_{0}\sin\theta\cos\phi}\\ \cdots\\ e^{j\frac{2\pi}{\lambda}x_{n}\sin\theta\cos\phi}\\ \cdots\\ e^{j\frac{2\pi}{\lambda}x_{N-1}\sin\theta\cos\phi}\end{bmatrix}$, (5)
@@
$\mathbf{a}_{z}(\theta)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}z_{0}\cos\theta}\cdots e^{j\frac{2\pi}{\lambda}z_{m}\cos\theta}\cdots e^{j\frac{2\pi}{\lambda}z_{M-1}\cos\theta}\end{bmatrix}^{H}.$ (6)
@@
$\tilde{\mathbf{a}}_{z}(\theta)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}x_{0}\sin\theta\cos(\phi+\Delta_{\phi})}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}x_{n}\sin\theta\cos(\phi+\Delta_{\phi})}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}x_{N-1}\sin\theta\cos(\phi+\Delta_{\phi})}\end{bmatrix}$, (16)
@@
$\tilde{\mathbf{a}}_{x}(\theta,\phi)=\begin{bmatrix}e^{j\frac{2\pi}{\lambda}z_{0}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}z_{m}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\\ \vdots\\ e^{j\frac{2\pi}{\lambda}z_{M-1}(\cos\theta\cos\Delta_{\theta}-\sin\theta\sin\Delta_{\theta}\sin(\phi+\Delta_{\phi}))}\end{bmatrix}$. (17)
@@ III Proposed Recursive Rotation Root-MUSIC (RR-Root-MUSIC) Method
Due to directive antenna patterns, the sensing power is significantly attenuated when the emitter direction deviates from the normal vector. Therefore, in this section, a RR-Root-MUSIC method is proposed. As shown in Fig. 3LABEL:sub@a, when the emitter direction is far away from the normal vector of array, Root-MUSIC method is applied at the initial orientation which provides a coarse estimation value at the lower directional gain. According to this estimation angle, the array is rotated into the main-lobe region as Fig. 3LABEL:sub@b, which increases the sensing power and improves the estimation condition. With the updated orientation, a new estimation angle is obtained, and the process repeats. As shown in Fig. 3LABEL:sub@c, the offset angle gradually shrinks over iterations, leading to higher effective SNR and eventual convergence toward the lower RMSE.
@@ IV Derived CRLB for Rotatable Array
The CRLB provides a benchmark for the minimum achievable RMSE of any unbiased estimator. According to literature [34], [35] and [36], to evaluate the performance of the proposed RR-Root-MUSIC method, the CRLB of the rotation array is derived. The derivation process is as follows.
@@
$\tilde{\mathbf{y}}=\big[\tilde{y}_{0,0},\,\tilde{y}_{0,1},\dots,\tilde{y}_{M-1,N-1}\big]^{T},$ (38)
@@
$\boldsymbol{\theta}=[\theta,\phi]^{T}.$ (39)
@@
$\mathbf{F}({\color[rgb]{0,0,1}\boldsymbol{\theta}})=\frac{2\,K\,s^{2}}{\sigma^{2}}\begin{bmatrix}P&Q\\ Q&R\end{bmatrix}$. (49)
@@
$CRLB=\mathbf{F}^{-1}(\boldsymbol{\theta})=\frac{\sigma^{2}}{2Ks^{2}}\frac{1}{PR-Q^{2}}\begin{bmatrix}R&-Q\\ -Q&P\end{bmatrix}$. (59)
Therefore, the lower bound of the RMSE of the $\hat{\theta}$ and $\hat{\phi}$ can be expressed as
@@ V Simulation Results
In this section, we present simulation results to evaluate the performance of proposed RR-Root-MUSIC method and the corresponding CRLB as a performance benchmark. The parameters of simulation are set as: $N=M=6$, $P_{t}$= 20dBm, $\sigma^{2}$= -100dBm, $\phi=90^{\circ}$, $r$ = 250m, $\lambda$ = 0.125, $A$ = $\lambda^{2}/4\pi$. In our simulation, all results are averaged over 2000 Monte Carlo realizations and RMSE is calculated to indicate the performance, which is given by
$\text{RMSE}=\sqrt{\frac{1}{L}\sum_{l}^{L}(\theta_{l}-\theta)^{2}},$ (68)
@@
In Fig. 4, the RMSE convergence curves of the proposed RR-Root-MUSIC method are illustrated for SNR = -10 dB and three different elevation angles: $\theta=75^{\circ}$, $\theta=45^{\circ}$, $\theta=15^{\circ}$, with CRLB as a performance benchmark, where $L=1$ in Fig. 4LABEL:sub@fig:conver-curve-low_L1 and $L=1000$ in in Fig. 4LABEL:sub@fig:conver-curve-low_L1000. Obviously, given a fixed value of SNR, the proposed method asymptotically converges to the corresponding error floor after about 13 or less rotations on average. As the angle of the emitter deviates away from the normal vector of the rotation array, the RMSE of the harvested performance gain ranges from one-order magnitude to three-order magnitude as $\theta$ varies from $75^{\circ}$ to $15^{\circ}$. This means that the larger the angle difference between the incident direction of the emitter and the normal direction of array, the larger the achieved performance gain by the proposed rotation operation. When $\theta$ is close to $90^{\circ}$, the proposed method takes only one time to converge to the same error floor. In summary, if the emitter direction is far away from the normal vector, a significant RMSE performance gain may be achieved.
@@
Figure 7: RMSE versus the elevation angles $\theta$ for proposed RR-Root-MUSIC method with transmit powers $P_{t}\in$ {0 dBm, 20 dBm}.
In Fig. 7, the RMSE curves of the proposed RR-Root-MUSIC method versus the elevation angle $\theta$ are shown for different transmitter powers $P_{t}\in$ {0 dBm, 20 dBm}, with CRLB as a performance benchmark. It is obvious that as $\theta\rightarrow 0^{o}$, the performance gain of the proposed RR-Root-MUSIC over conventional fixed Root-MUSIC grows gradually. In extreme situation, i.e., $\theta\leq 5^{o}$, the RMSE performance of the proposed method is more four-order magnitude better than that of the conventional fixed Root-MUSIC. This further confirms that a rotatable array may effectively mitigate the pattern-induced performance loss.
@@
(b) RMSE versus the motion path of UAV
Figure 8: The motion path of the UAV around BS, and RMSE of proposed RR-Root-MUSIC method versus the motion path of UAV.
To observe the influence of the proposed rotation method in low-altitude UAV flying, Fig. 8LABEL:sub@fig:uav_path shows the motion path of the UAV around BS. Fig. 8LABEL:sub@fig:MSE_uav illustrates the corresponding RMSE performance curves along semi-circular trajectory of UAV for the proposed RR-Root-MUSIC and the traditional fixed MUSIC method. From Fig. 8LABEL:sub@fig:uav_path, the UAV flies form (0, -250, 0) and (0, -150, 0) to (0, 250, 0) and (0, 150, 0) with radii $r=250$ and $r=150$, respectively. Compared with the traditional Root-MUSIC method, the proposed RR-Root-MUSIC method makes a dramatic performance gain over the conventional fixed Root-MUSIC when the position of the UAV is on top of BS, which is usually a blind coverage area. Additionally, due to the use of rotatable array, the sensing performance keeps stable during the flight process of UAV. Thus, the proposed rotatable array system and method will offer an excellent new solution to the future low-latitude communication network.
@@ VI Conclusions
In this paper, to solve the low-altitude coverage problem of MIMO transceiver with directional antenna elements in practice, a rotatable array system was established with each element being the same directional pattern. Here, via a sensing and rotation iterative loop, the RR-Root-MUSIC was proposed to fully harvest the antenna array gain to make a dramatic performance enhancement, especially for the scenario that the emitter direction is far away from the normal vector of array. Compared with classical Root-MUSIC without rotation, the proposed method may achieve up to -order magnitude improvement or more at the elevation angle $\theta\leq 15^{o}$ in terms of RMSE performance. Moreover, the proposed RR-Root-MUSIC converges fast and provides a full coverage on such an area on top of BS, which cannot be covered by the conventional fixed antenna array system like 5G or 4G. Due to the above advantages of rotation-array system, it will become an extremely potential technique for the future low-altitude wireless network to completely break the bottleneck of low-latitude coverage faced by the current mobile communication like 5G.